OpenStudy (anonymous):

NEED HELP! What are the intervals of increase and decrease of f(x) = x ( sqrt of x+3) * This is a calc problem

OpenStudy (freckles):

Have you considered finding the derivative?

OpenStudy (usukidoll):

we need the first derivative test... wait we need to find critical points first!

OpenStudy (freckles):

f'>0 tells us f is increasing f'<0 tells us f is decreasing Find the critical numbers and consider the domain.

OpenStudy (usukidoll):

what @freckles said is correct... use product rule

OpenStudy (anonymous):

I already did all that, and I got (sqrt x+3) + x/(2(sqrt x +3)). But I just dont know how to go about solving for the critical values

OpenStudy (freckles):

I don't think we need to find the x-intercepts to find where the function is increasing/decreasing.

OpenStudy (anonymous):

is it just x= -3?

OpenStudy (usukidoll):

noooooo....................... freckles we need the critical points first or we can't go further T_T

OpenStudy (usukidoll):

sorry mary that comment was at freckles

OpenStudy (freckles):

\[f'(x)=\sqrt{x+3}+\frac{x}{2 \sqrt{x+3}} \\ \text{ solve } f'(x)=0 \text{ \to find the critical numbers }\]

OpenStudy (zale101):

To get the critical points, we first need to set the derivative to zero

OpenStudy (anonymous):

Yea I did that already guys!!!

OpenStudy (zale101):

and solve for x. Those x's are the crtical points.

OpenStudy (freckles):

good then what did you for the x's?

OpenStudy (anonymous):

I just need help in solving for x

OpenStudy (freckles):

\[\sqrt{x+3} \frac{2\sqrt{x+3}}{2\sqrt{x+3}}+\frac{x}{2 \sqrt{x+3}}\] find common denominator

OpenStudy (anonymous):

No, we find the first derivative and then find the critical points

OpenStudy (freckles):

\[\frac{2 (x+3)+x}{2 \sqrt{x+3}}=0\]

OpenStudy (anonymous):

hey freckles, how did you get +x in the numerator?

OpenStudy (freckles):

(sqrt x+3) + x/(2(sqrt x +3) isn't this what you got for f'?

OpenStudy (anonymous):

OH, K i see it

OpenStudy (freckles):

\[f'(x)=\sqrt{x+3}+\frac{x}{ 2 \sqrt{x+3}} \\ f'(x) =\frac{ \sqrt{x+3} }{1} \cdot \frac{2 \sqrt{x+3}}{2 \sqrt{x+3}} +\frac{x}{2 \sqrt{x+3}} \\ f'(x)=\frac{2(\sqrt{x+3})^2+x}{2 \sqrt{x+3}}\]

OpenStudy (anonymous):

So is it x=-2 for when f'(x) = 0

OpenStudy (freckles):

yes

OpenStudy (anonymous):

and then the denominator would give me when derivative is undefined?

OpenStudy (anonymous):

so the critical points are x= -2, and -3?

OpenStudy (freckles):

in by the way the domain of f is x>=-3 so we have this so far: